# A Langevin sampling algorithm inspired by the Adam optimizer

### v1

@@ Abstract
We present a framework for adaptive-stepsize MCMC sampling based on time-rescaled Langevin dynamics, in which the stepsize variation is dynamically driven by an additional degree of freedom. Our approach augments the phase space by an additional variable which in turn defines a time reparameterization. The use of an auxiliary relaxation equation allows accumulation of a moving average of a local monitor function and provides for precise control of the timestep while circumventing the need to modify the drift term in the physical system. Our algorithm is straightforward to implement and can be readily combined with any off-the-peg fixed-stepsize Langevin integrator. As a particular example, we consider control of the stepsize by monitoring the norm of the log-posterior gradient, which takes inspiration from the Adam optimizer, the stepsize being automatically reduced in regions of steep change of the log posterior and increased on plateaus, improving numerical stability and convergence speed. As in Adam, the stepsize variation depends on the recent history of the gradient norm, which enhances stability and improves accuracy compared to more immediate control approaches. We demonstrate the potential benefit of this method–both in accuracy and in stability–in numerical experiments including Neal’s funnel and a Bayesian neural network for classification of MNIST data.
@@ 1 Introduction
Monte Carlo sampling schemes are ubiquitous in modern-day science, engineering, and finance. They are used to quantify risk and uncertainty, parameterize statistical models, calculate thermodynamic quantities in physical models, and explore protein configurational states. The aim in sampling is to generate independent and identically distributed (i.i.d.) realizations $x_{i}$ of a random variable $X\in\mathbb{R}^{d}$ distributed according to a given probability law, which we assume is defined in terms of a positive, smooth density $\pi$. Such samples can then be used to estimate probabilities of certain events or expectations of functions, to calculate uncertainties, assess or compare models, or to explore optimality of parameterizations in machine learning applications. One of the most powerful and widely used categories of sampling methods are Markov Chain Monte Carlo schemes (MCMC [3, 13]), which generate the samples using Markov Chains that are ergodic with respect to the target probability measure. The original MCMC scheme is the Metropolis-Hastings (MH) algorithm [66, 40] which uses random proposals in conjunction with an accept-reject procedure (the MH-criterion) to generate the Markov Chain. The MH framework is very general and allows many alternatives for proposal generation. The use of Metropolis correction may, however, add substantial computational burden and the rejection steps can slow convergence to the target distribution.
@@
One approach to tackle some of the mentioned issues comes in the form of adaptive stepsize methods, which are widely used in integration of deterministic systems, e.g., $\frac{\mathop{}\!\text{d}}{dt}x=\phi(x)$. Variable stepsize procedures try to estimate the local discretization error being introduced at each step and adjust the stepsize down or up to maintain a certain prescribed local error tolerance (see, e.g., Chapter 3 of [54]). The error estimate may be based on finite difference approximation, or extrapolation, or the use of specialized ‘embedded’ integration schemes (e.g., Runge-Kutta Fehlberg methods [36]). These techniques have also been adapted for weak approximation (sampling) using stochastic differential equations [84, 97, 94]. Implementations of this approach are in widespread use in modern simulation software. The molecular dynamics package OpenMM [35] uses a variable stepsize scheme in which local errors are estimated on-the-fly via simple Euler discretization (or Euler-Maruyama discretization for their LD integrator; see [35]). Additionally, there has been recent developments of adaptive stepsize methods, which allow for high strong order approximation in [47] and [37]. We also mention the development of recent methodology for local stepsize adjustment for HMC based on Gibbs self tuning in [9] and local adaptation of other parameters in HMC [10].
@@
$\frac{\mathop{}\!\text{d}}{\mathop{}\!\text{d}\tau}x=R(x)\phi(x),\quad\frac{\mathop{}\!\text{d}t}{\mathop{}\!\text{d}\tau}=R(x),$ (4)
@@
By combining the methodology of time-transformed SDEs with Adam’s approach to adjust the stepsize based on a moving average over recent history, we derive a flexible framework for adaptive-stepsize sampling, which we call “SamAdams” (sampling with adaptive moderated stepsize). The procedure can be combined with state-of-the art integrators for Langevin dynamics. In particular, it is possible to transform the splitting integrators mentioned above into adaptive algorithms. Adaptivity improves numerical stability compared to relying on a constant stepsize, as demonstrated in Fig. 1 where a fixed stepsize Langevin integrator is compared with its adaptive counterpart.
@@ 2 Sundman-transformed SDEs and Averaging
In [61], it was shown that the canonical distribution $\rho_{\beta}$ is no longer invariant under the process, so that work introduced an additional drift correction term to recover useful samples.
@@
with hyperparameter $\alpha>0$ representing the attack rate of a relaxation process. The monitor function $g(x,p)$ can be any positive smooth function and ultimately decides on what basis a small time increment $\mathop{}\!\text{d}t$ is varied. The rescaling function $\psi$, over a crucial interval, mimics a reciprocal power of $\zeta$. If the monitor function takes on larger values, $\zeta$ will tend to increase and the time interval $\mathop{}\!\text{d}t$ will shrink (implying smaller stepsizes during numerical integration). If the monitor function is small (meaning the solution is locally smooth and easy to integrate), the rescaling will lead to an increase in stepsize. This corresponds to the paradigm to take as large a stepsize as possible but as small as necessary. The dynamics (9) effectively computes a moving average of the driving function $g$ over the recent history, exponentially weighted with rate $\alpha$ (see the discussion in Appendix C). Making the stepsize adaptation depend on the recent history of a driving function is one reason why the Adam optimizer became so successful.
Since one can employ many different monitor functions $g(x,p)$ and Sundman transform kernels $\psi(\zeta)$, we obtain a flexible framework for adaptive stepsize sampling, given by [^1] A version of this system for generic SDEs is given in the Supplementary Material.
@@
In Sec. 4 we show that we can pick the Sundman transform $\psi$ and driving function $g$ to obtain a sampler that resembles the Adam optimizer, although it is important to emphasize that the framework is very general. We next address the ergodicity of process (10)-(13), how to obtain canonical averages from it, and how to simulate it in practice.
@@ 2.1 Ergodic Averages
where the second and third lines follow from ergodicity and the fourth line from identifying $\psi\mathop{}\!\text{d}\tau=\mathop{}\!\text{d}t$. We remark that the process (10)-(13) can be shown to be ergodic under minimal assumptions as in [85] which treats a more general case. Similar techniques can be found in [63, 61, 76]. In particular, under sufficient smoothness, the assumption that the time-rescaling $\psi$ is uniformly bounded from below and above and the force is convex outside a ball, one can show ergodicity of the time-rescaled process (10)-(13) by considering the family of Lyapunov functions considered in [63]. Resulting in an ergodicity result with a convergence rate which depends on the uniform lower bound on $\psi$. More sophisticated techniques would be required to show an improved convergence rate of the Adam sampler, akin to the adaptive stepsize results available in the optimization literature (see, for example, [27] for Adam or Adagrad).
@@
$\frac{\sum_{i=1}^{N}\phi(x^{i}_{\tau},p^{i}_{\tau})\psi(\zeta^{i}_{\tau})}{\sum_{i=1}^{N}\psi(\zeta^{i}_{\tau})}\approx\frac{\mathbb{E}_{\Pi_{\tau}}(\phi R)}{\mathbb{E}_{\Pi_{\tau}}(R)}=\mathbb{E}_{\pi}(\phi).$ (15)
@@
The subscript $i$ in the upper right-hand corner of the diagram can either denote different points in time along a single trajectory (time-average) or a trajectory index when averaging over samples of different trajectories taken at the same time (trajectory-average).
@@ 3 Numerical Integration
In order to simulate the time rescaled dynamics (10) to (13), we need to discretize the continuous-time process with a suitable numerical integrator. For convenience we adopt the framework of [91, 16, 11, 55] in which a symplectic splitting of the Hamiltonian part of the underdamped Langevin system is composed with a map that exactly preserves the momentum distribution. A frequent choice for the stochastic part is
@@ 3.1 The Algorithm
A procedure for implementation of a symmetric variant of SamAdams is given in Alg. 1.
Algorithm 1 SamAdams
Given: $\hat{\Phi}_{\Delta t}$ a (fixed-stepsize) Langevin integrator, ${\psi}$ a suitable Sundman transformation.
@@
     $({x}_{n+1},{p}_{n+1}):=\hat{\Phi}_{\Delta t_{n+1}}(x_{n},p_{n})$. $\triangleright$ Langevin dynamics step with stepsize $\Delta t_{n+1}$.
@@
We can name integrators from the SamAdams family simply by introducing letters Z corresponding to updates of the $\zeta$ dynamics. Palindromic letter sequences like ZBAOABZ indicate a symmetric introduction of $\zeta$ updates around a fixed stepsize method. Compared to the alternative ZBAOAB, the symmetric inclusion of the two Z half-steps has a very significant consequence, the effect of which can be seen in experiments: the weights $\mu_{n}$ that are output with the samples are more accurate than if we for example relied on the stepsize that was used to advance the previous time-step.
In this method the implementation of the underlying Langevin integrator itself is as in the fixed-stepsize setting. There is no additional significant cost overhead above the usual cost of fixed stepsize integration if, for example, the BAOAB integrator is used for Langevin dynamics, since the force evaluation needed to update Z is already performed in the BAOAB iteration.
We note that initialization of Algorithm 1 requires a choice for $\zeta_{0}$. In our experiments we typically chose the initial conditions $(x,p)$ so that the force was small and thus used an intial value of $\zeta_{0}=0$. Since $\psi(\zeta_{0})=M$ for either of the given choices of the filter function, we have $\Delta t_{0}=M\Delta\tau$. It is expected that there will be some equilibration of the timestep control in the first few steps of integration.
@@ 3.2 Convergence and Order of Accuracy
The order of accuracy of a numerical method for SDEs can be studied in different contexts. In some disciplines, the quantity of interest is the strong accuracy, defined as the fidelity of the numerical solution to the solution of the SDE associated to some realization of the Wiener process. For sampling, the more relevant quantity is the accuracy of the approximation of the distribution generated by SDE solutions at a specified time (weak accuracy). An article of G. Vilmart [99, Proposition 6.1] provides a rigorous proof of the weak convergence of general splitting methods, which involve components which can be integrated exactly (in the weak sense) but which may contain multiplicative noise, as is the case for ZBAOABZ and other such integrators for SamAdams. Specifically we can state a theorem regarding finite time approximation based on this work (see Theorem 1).
@@
Consider the system (10)-(13) and assume that $\psi$, $\sqrt{\psi}$, $\nabla U$, $g$ are $C^{6}$ functions with all partial derivatives bounded, further assume that $\psi>r$ for some $r>0$. Then consider splitting into A, B, O and Z components and generating a sequence of points $(x_{n},p_{n},\mu_{n})$ based on Algorithm 1 with any symmetric splitting with these components. Let $\chi(\cdot)\in C^{6}$ be an observable function of $x,p$ where all partial derivatives have polynomial growth, then for all $\Delta\tau k\leq T$
@@
We first remark that $\mathbb{E}(\mu_{n})>0$ and $\mathbb{E}(\psi(\zeta(n\Delta\tau)))>0$ for all $n\in\mathbb{N}$ due to the uniform bound assumption on $\psi$. Let $x=\mathbb{E}(\chi(x_{n},p_{n})\mu_{n})$, $y=\mathbb{E}(\mu_{n})$, $w=\mathbb{E}(\chi(X(n\Delta\tau),P(n\Delta\tau))\psi(\zeta(n\Delta\tau))$ and $z=\mathbb{E}(\psi(\zeta(n\Delta\tau))$. Then we have
@@
and we have that $|x-w|\leq C\Delta\tau^{2}$ and $|z-y|\leq C\Delta\tau^{2}$ by [99, Proposition 6.1], then due to the uniform lower bound on $\psi$ we have the required result. ∎
@@
If initialized according to the invariant measure of (10)-(13) and if the process is ergodic we have that
$\frac{\mathbb{E}(\chi(X(n\Delta\tau),P(n\Delta\tau))\psi(\zeta(n\Delta\tau))}{\mathbb{E}(\psi(\zeta(n\Delta\tau))}=\mathbb{E}_{\pi_{\beta}}(\chi)$
for all $n\in\mathbb{N}$ following from Section 2.1.
The weak convergence can in principle be studied in the asymptotic sense as the time interval tends to infinity, i.e., we may consider the asymptotic evolution of the weak error. The challenge then is to establish the convergence rate in a suitable framework and to study the systematic bias that is introduced due to discretization. Theoretical study of the geometric convergence of Alg. 1 and the order of accuracy of the stationary distribution of the numerical method will be explored in future work, using techniques that are by now well developed in the setting of Langevin dynamics [11, 57, 34, 23, 68, 60].

### v2

@@ Abstract
We present a framework for adaptive-stepsize MCMC sampling based on time-rescaled Langevin dynamics, in which the stepsize variation is dynamically driven by an additional degree of freedom. The use of an auxiliary relaxation equation allows accumulation of a moving average of a local monitor function and provides for precise control of the timestep while circumventing the need to modify the drift term in the physical system. Our algorithm is straightforward to implement and can be readily combined with any off-the-peg fixed-stepsize Langevin integrator. As a particular example, we consider control of the stepsize by monitoring the norm of the log-posterior gradient, which takes inspiration from the Adam optimizer, the stepsize being automatically reduced in regions of steep change of the log posterior and increased on plateaus, improving numerical stability and convergence speed. As in Adam, the stepsize variation depends on the recent history of the monitor function, which enhances stability and improves accuracy compared to more immediate control approaches. We demonstrate the potential benefit of this method–both in accuracy and in stability–in numerical experiments including Neal’s funnel and a Bayesian neural network for classification of MNIST data.
@@ 1 Introduction
Monte Carlo sampling schemes are ubiquitous in modern-day science, engineering, and finance. They are used to quantify risk and uncertainty, parameterize statistical models, calculate thermodynamic quantities in physical models, and explore protein configurational states. The aim in sampling is to generate independent and identically distributed (i.i.d.) realizations $x_{i}$ of a random variable $X\in\mathbb{R}^{d}$ distributed according to a given probability law, which we assume is defined in terms of a positive, smooth density $\pi$. Such samples can then be used to estimate probabilities of certain events or expectations of functions, to calculate uncertainties, assess or compare models, or to explore optimality of parameterizations in machine learning applications. One of the most powerful and widely used categories of sampling methods are Markov Chain Monte Carlo schemes (MCMC [3, 13]), which generate the samples using Markov Chains that are ergodic with respect to the target probability measure. The original MCMC scheme is the Metropolis-Hastings (MH) algorithm [68, 41] which uses random proposals in conjunction with an accept-reject procedure (the MH-criterion) to generate the Markov Chain. The MH framework is very general and allows many alternatives for proposal generation. The use of Metropolis correction may, however, add substantial computational burden and the rejection steps can slow convergence to the target distribution. Even with a warm-start assumption the convergence rate is known to scale polynomially with respect to dimension [85, 106].
@@
One approach to tackle some of the mentioned issues comes in the form of adaptive stepsize methods, which are widely used in integration of deterministic systems, e.g., $\frac{\mathop{}\!\text{d}}{dt}x=\phi(x)$. Variable stepsize procedures try to estimate the local discretization error being introduced at each step and adjust the stepsize down or up to maintain a certain prescribed local error tolerance (see, e.g., Chapter 3 of [55]). The error estimate may be based on finite difference approximation, or extrapolation, or the use of specialized ‘embedded’ integration schemes (e.g., Runge-Kutta Fehlberg methods [37]). These techniques have also been adapted for weak approximation (sampling) using stochastic differential equations [87, 101, 98]. Implementations of this approach are in widespread use in modern simulation software. The molecular dynamics package OpenMM [36] uses a variable stepsize scheme in which local errors are estimated on-the-fly via simple Euler discretization (or Euler-Maruyama discretization for their LD integrator; see [36]). Additionally, there has been recent developments of adaptive stepsize methods, which allow for high strong order approximation in [48] and [38]. We also mention the development of recent methodology for local adaptation of stepsize and other parameters within Metropolis-based algorithms, for example, HMC and more general PDMP-based algorithms (see [9, 10, 22]).
@@
$\frac{\mathop{}\!\text{d}x}{\mathop{}\!\text{d}\tau}=R(x)\phi(x),\quad\frac{\mathop{}\!\text{d}t}{\mathop{}\!\text{d}\tau}=R(x),$ (4)
@@
By combining the methodology of time-transformed SDEs with Adam’s approach to adjust the stepsize based on a moving average over recent history, we derive a flexible framework for adaptive-stepsize sampling, which we call ”SamAdams” (sampling with adaptive moderated stepsize). The procedure can be combined with state-of-the art integrators for Langevin dynamics. In particular, it is possible to transform the splitting integrators mentioned above into adaptive algorithms. Adaptivity improves numerical stability compared to relying on a constant stepsize, as demonstrated in Fig. 1 where a fixed stepsize Langevin integrator is compared with its adaptive counterpart.
@@ 2 Sundman-transformed SDEs and Averaging
In [62], it was shown that the canonical distribution $\rho_{\beta}$ is no longer invariant under the process. To correct the measure that article introduced a drift correction
$\Delta(x):=\frac{\gamma}{\beta}\nabla R(x).$
This method adds some complexity in the form of an additional gradient computation.
@@
with hyperparameter $\alpha>0$ representing the attack rate of a relaxation process. The monitor or driving function $g(x,p)$ can be any positive smooth function and ultimately decides on what basis a small time increment $\mathop{}\!\text{d}t$ is varied. The rescaling function $\psi$, over a crucial interval, mimics a reciprocal power of $\zeta$. If the monitor function takes on larger values, $\zeta$ will tend to increase and the time interval $\mathop{}\!\text{d}t$ will shrink (implying smaller stepsizes during numerical integration). If the monitor function is small (meaning the solution is locally smooth and easy to integrate), the rescaling will lead to an increase in stepsize. This corresponds to the paradigm to take as large a stepsize as possible but as small as necessary. Making the stepsize adaptation depend on the recent history of a driving function is one reason why the Adam optimizer became so successful.
Since one can employ many different monitor functions $g(x,p)$ and Sundman transform kernels $\psi(\zeta)$, we obtain a flexible framework for adaptive stepsize sampling, given by [^1] A version of this system for generic SDEs is given in Appendix A.
@@
Note that in the Adam optimizer, the stepsizes are adapted based on moving averages (see Appendix B for a detailed discussion). The dynamics (12) also effectively computes a moving average of the driving function $g$ over the recent history, exponentially weighted with rate $\alpha$, which is then used to modify the stepsize $\Delta t$ (see also Appendix C). Due to the stepsize adaptation through moving averages and the fact that, as we will show in Sec. 4, one can design transform kernel $\psi$ and driving function $g$ in a way to resemble the Adam equations, the dynamics (10)-(13) can be understood as an Adam-inspired sampling framework. We name this framework SamAdams (sampling with adaptively moderated stepsizes). Although we will focus on particular choices for $\psi$ and $g$, it is important to emphasize that the framework is very general. We next address the ergodicity of process (10)-(13), how to obtain canonical averages from it, and how to simulate it in practice.
@@ 2.1 Ergodic Averages
where the second and third lines follow from ergodicity and the fourth line from identifying $\psi\mathop{}\!\text{d}\tau=\mathop{}\!\text{d}t$. We remark that the process (10)-(13) can be shown to be ergodic under minimal assumptions as in [88] which treats a more general case. Similar techniques can be found in [64, 78]. In particular, under sufficient smoothness, the assumption that the time-rescaling $\psi$ is uniformly bounded from below and above and the force is convex outside a ball, one can show ergodicity of the time-rescaled process (10)-(13) by considering the family of Lyapunov functions considered in [64]. Resulting in an ergodicity result with a convergence rate which depends on the uniform lower bound on $\psi$. More sophisticated techniques would be required to show an improved convergence rate of the Adam sampler, akin to the adaptive stepsize results available in the optimization literature (see, for example, [28] for Adam or Adagrad).
@@
$\frac{\sum_{i=1}^{N}\phi(x^{i}_{\tau},p^{i}_{\tau})\psi(\zeta^{i}_{\tau})}{\sum_{i=1}^{N}\psi(\zeta^{i}_{\tau})}\approx\frac{\mathbb{E}_{\Pi_{\tau}}(\phi\psi)}{\mathbb{E}_{\Pi_{\tau}}(\psi)}=\mathbb{E}_{\pi}(\phi).$ (15)
@@
The subscript $n$ in the upper right-hand corner of the diagram can either denote different points in time along a single trajectory (time-average) or a trajectory index when averaging over samples of different trajectories taken at the same time (trajectory-average).
@@ 3 Numerical Integration
In order to simulate SamAdams, we need to discretize the continuous-time process (10) to (13) with a suitable numerical integrator. For convenience we adopt the framework of [94, 16, 11, 56] in which a symplectic splitting of the Hamiltonian part of the underdamped Langevin system is composed with a map that exactly preserves the momentum distribution. A frequent choice for the stochastic part is
@@ 3.1 The Algorithm
Using the splitting components from the previous section, we can write down algorithms to simulate SamAdams. With slight notational ambiguity, we use integer subscripts to denote timesteps in artificial time, thus $x_{n}\approx x_{\tau}(\tau_{n})$, etc. Algorithm 1 presents the ZBAOABZ method that we have used in all our numerical experiments.
Algorithm 1 SamAdams (ZBAOABZ)
Given: $\hat{\Phi}^{\text{BAOAB}}_{\Delta t}$ the (fixed-stepsize) BAOAB integrator, ${\psi}$ a suitable Sundman transformation.
@@
     $({x}_{n+1},{p}_{n+1}):=\hat{\Phi}^{\text{BAOAB}}_{\Delta t_{n+1}}(x_{n},p_{n})$. $\triangleright$ BAOAB step with stepsize $\Delta t_{n+1}$.
@@
The implementation of the underlying Langevin integrator, here BAOAB (denoted by $\hat{\Phi}^{\text{BAOAB}}_{\Delta t}$ in Alg. 1), is as in the fixed-stepsize setting. There is no additional significant cost overhead above the usual cost of fixed stepsize integration, since the force evaluation needed to update Z is already performed in the BAOAB iteration. However, one could use other fixed-stepsize Langevin dynamics integrators as well.
Palindromic letter sequences like ZBAOABZ indicate a symmetric introduction of $\zeta$ updates around the fixed-stepsize method BAOAB. Compared to the alternative ZBAOAB, the symmetric inclusion of the two Z-half-steps has a very significant consequence, the effect of which can be seen in experiments: the weights $\mu_{n}$ that are output with the samples are more accurate than if we for example relied on the stepsize that was used to advance the previous time-step, see Appendix D.
We note that initialization of Algorithm 1 requires a choice for $\zeta_{0}$. There are different approaches to this depending on the particular choices for $\psi$ and $g$, and we discuss reasonable options in Sec. 5. Independently of $\zeta_{0}$, it is expected that there will be some equilibration of the timestep control in the first few steps of integration.
A generic version of Algorithm 1 for overdamped (first order) dynamics is given Appendix A.
@@ 3.2 Convergence and Order of Accuracy
The order of accuracy of a numerical method for SDEs can be studied in different contexts. In some disciplines, the quantity of interest is the strong accuracy, defined as the fidelity of the numerical solution to the solution of the SDE associated to some realization of the Wiener process. For sampling, the more relevant quantity is the accuracy of the approximation of the distribution generated by SDE solutions at a specified time (weak accuracy). An article of G. Vilmart [103, Proposition 6.1] provides a rigorous proof of the weak convergence of general splitting methods, which involve components which can be integrated exactly (in the weak sense) but which may contain multiplicative noise, as is the case for ZBAOABZ and other such integrators for SamAdams. Specifically, we can state a theorem regarding finite time approximation based on this work (see Theorem 1).
@@
Consider the system (10)-(13) and assume that $\psi$, $\sqrt{\psi}$, $\nabla U$, $g$ are $C^{6}$ functions with all partial derivatives bounded, further assume that $r<\psi<M$ for some $r,M>0$. Then consider a splitting of the form ZBAOABZ and generating a sequence of points $(x_{n},p_{n},\mu_{n})$ based on Algorithm 1. Let $\chi(\cdot)\in C^{6}$ be an observable function of $x,p$ where all partial derivatives have polynomial growth, then for all $\Delta\tau k\leq T$
@@
We first remark that $\mathbb{E}(\mu_{n})>r$ and $\mathbb{E}(\psi(\zeta(n\Delta\tau)))>r$ for all $n\in\mathbb{N}$ due to the uniform bound assumption on $\psi$. Let $x=\mathbb{E}(\chi(x_{n},p_{n})\mu_{n})$, $y=\mathbb{E}(\mu_{n})$, $w=\mathbb{E}(\chi(X(n\Delta\tau),P(n\Delta\tau))\psi(\zeta(n\Delta\tau))$ and $z=\mathbb{E}(\psi(\zeta(n\Delta\tau))$. Then we have
@@
and we have that $|x-w|\leq C\Delta\tau^{2}$ and $|z-y|\leq C\Delta\tau^{2}$ by [103, Proposition 6.1] applied to the Strang splitting between $Z$ and the symmetric kinetic Langevin dynamics integrator (for example BAOAB), then due to the uniform lower bound on $\psi$ we have the required result. ∎
@@
Theorem 1 also applies to for example ZUBUZ (or replacing the middle with any other weak order two integrator for underdamped Langevin dynamics, for example, OBABO or ABOBA).
Remark 2.
If, in addition, the process (10)-(13) is ergodic, then
$\lim_{n\to\infty}\frac{\mathbb{E}(\chi(X(n\Delta\tau),P(n\Delta\tau))\psi(\zeta(n\Delta\tau))}{\mathbb{E}(\psi(\zeta(n\Delta\tau))}=\mathbb{E}_{\pi_{\beta}}(\chi)$
following from Section 2.1.
The weak convergence can in principle be studied in the asymptotic sense as the time interval tends to infinity, i.e., we may consider the asymptotic evolution of the weak error. The challenge then is to establish the convergence rate in a suitable framework and to study the systematic bias that is introduced due to discretization. Theoretical study of the geometric convergence of Alg. 1 and the order of accuracy of the stationary distribution of the numerical method will be explored in future work, using techniques that are by now well developed in the setting of Langevin dynamics [11, 58, 35, 24, 70, 61]. In Appendix E we examine the order of accuracy numerically.